OpenStudy (anonymous):

Suppose f(x) is continuous on [a; b] and is revolved about the x-axis. The surface area of the resulting fi gure is given by integral from a to be of 2(pi)y * sqrt(1+(dy/dx)^2) -Consider the function f(x) = 1/x on the interval [1; infinity). Show that the surface area of f revolved about the x-axis on the same interval is infinite.

OpenStudy (turingtest):

have you set up the integral yet?

OpenStudy (anonymous):

\[\int\limits_{1}^{\infty}\frac{ 1 }{ x }dx\] this one?

OpenStudy (turingtest):

what happened to the sqrt term?

OpenStudy (anonymous):

you mean this one? \[\int\limits_{a}^{b}2 \pi y \sqrt{1+(\frac{ dy }{ dx }})^2\]

OpenStudy (turingtest):

yeah, but use y=1/x..

OpenStudy (anonymous):

so does it become \[\int\limits_{1}^{\infty}2 \pi (\frac{ 1 }{ x })\sqrt{1+(\frac{ -1 }{ x^2 }})^2\]

OpenStudy (turingtest):

yes, now my advice is to simplify it

OpenStudy (anonymous):

so: \[2 \pi \int\limits_{1}^{\infty}\frac{ 1 }{ x }\sqrt{1+(\frac{ 1 }{ x^4 }})\]

OpenStudy (turingtest):

right, now I'm not so sure of the best route from here, but it seems to me a good idea to pull 1/x^4 out of the radical

OpenStudy (turingtest):

you should really be writing dx, it is important at this point in calc

OpenStudy (anonymous):

oh yes ill remember to add dx! also should i try finding the limit as say "t" approaches infinity and make the integral go from 1 to "t"?

OpenStudy (turingtest):

yeah, but you needn't write that until you are about to evaluate. for now it seems to me that you want to do this with a somewhat tricky trig substitution. If there is a better way I don't see it yet.

OpenStudy (turingtest):

but first I would factor 1/x^4 out of the radical as I said

OpenStudy (anonymous):

so would it then become \[1/x^4\sqrt{x^4+1}\]

OpenStudy (anonymous):

then all together \[2 \pi \int\limits_{1}^{\infty}1/x^5\sqrt{x^4+1}\]

OpenStudy (turingtest):

you factored that wrong...

OpenStudy (turingtest):

\[\sqrt{1+\frac1{x^4}}=\sqrt{\frac {x^4+1}{x^4}}=\frac1{x^2}\sqrt{x^4+1}\]

OpenStudy (anonymous):

ah i see..

OpenStudy (anonymous):

Do you know about hyperbolic functions? Because I doubt that these kind of substitutions can be avoided here.

OpenStudy (turingtest):

I am thinking that what might be better is to do this along y instead

OpenStudy (anonymous):

If it can be simplified by changing a different method it's even better, the above integral is pretty complex (-:

OpenStudy (turingtest):

let \(x=1/y\), then as \(x\to\infty,~~y\to0\)\[S=2\pi\int_1^0 y\sqrt{1+\frac1{y^4}}dy\]

OpenStudy (turingtest):

not sure though, just a thought

OpenStudy (turingtest):

still a nasty integral...

OpenStudy (anonymous):

I thought I should write down some of the imperative steps to solve this integral, for completeness reasons of this post, however I would like the opener @peggiepenguin to make sure that we're talking about the same problem here. However: \[\Large \int \frac{1}{x^3}\sqrt{x^4+1}dx \] Setting \(x^2=\sinh \alpha\) such that \(dx/d\alpha=\cosh\alpha/(2\sqrt{\sinh \alpha}) \) Substituting this back yields to: \[\Large \int\frac{1}{\sqrt{\sinh^3\alpha}} \cdot\frac{\sqrt{\sinh^2\alpha+1}}{2\sqrt{\sinh\alpha}} \cdot\cosh\alpha d\alpha\] Hyperbolic trigonometry: \[\Large \cosh^2\alpha-\sinh^2\alpha=1 \longrightarrow \cosh^2\alpha=\sinh^2\alpha+1 \] Cleaning this all up: \[\Large \frac{1}{2}\int \frac{\cosh^2\alpha}{\sqrt{\sinh^4}\alpha}d\alpha=\frac{1}{2} \int\frac{\cosh^2\alpha}{|\sinh^2\alpha|}d\alpha\] Cheating a little, considering the numerator as positive within the domain of the integral: \[\Large \frac{1}{2}\int\frac{\cosh^2\alpha}{\sinh^2\alpha}d\alpha=\frac{1}{2}\int\coth^2\alpha d\alpha =\frac{1}{2}(\alpha-\coth\alpha)\]

OpenStudy (anonymous):

Trigonometric back-substitution: \[\Large \sinh\alpha=x^2=\frac{x^2}{1} \] and remember the hyperbolic Pythagoras \[\Large \cosh^2\alpha-\sinh^2\alpha=1 \\ \Large x^2-y^2=r^2 \] So: